# Parallel Key-Value Cache Fusion for Position Invariant RAG

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) underscore the necessity of Retrieval Augmented Generation (RAG) to leverage external information. However, LLMs are sensitive to the position of relevant information within contexts and tend to generate incorrect responses when such information is placed in the middle, known as ‘Lost in the Middle’ phenomenon. In this paper, we introduce a framework that generates consistent outputs for decoder-only models, irrespective of the input context order. Experimental results for three open domain question answering tasks demonstrate position invariance, where the model is not sensitive to input context order, and superior robustness to irrelevent passages compared to prevailing approaches for RAG pipelines.
@@
Philhoon Oh  Jinwoo Shin  James Thorne
KAIST AI {philhoonoh,jinwoos,thorne}@kaist.ac.kr
@@ 1 Introduction
In Retrieval Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2021; Izacard et al., 2022), models first extract relevant information from a knowledge base and then incorporate this extracted information with its parameteric knowledge to generate the response. This two-step approach is the de-facto approach for knowledge-intensive tasks (Lewis et al., 2021; Petroni et al., 2021).
However, decoder-only models exhibit an intrinsic positional bias, assigning more attention to tokens at the beginning or end of the input sequence while often overlooking relevant context located in the middle, a problem known as the ‘Lost in the Middle’ (Liu et al., 2023). Previous works to address this issue involves training with specific prompt (He et al., 2024) or data-intensive training (An et al., 2024). Other works aimed at modifying positional embeddings (Hsieh et al., 2024b) or reducing positional attention bias in LLMs (Yu et al., 2024a). Yet, none of these methods fully guarantee a solution to this intrinsic bias in LLMs for RAG.
@@
In this paper, we introduce a framework for decoder-only models, called Key Value Fusion (KV Fusion), to generate consistent outcomes regardless of input order as illustrated in Figure 1. KV Fusion consists of two components: a prefill decoder that extract key-values caches in parallel and a trainable decoder that utilizes extracted key value caches to produce consistent outcome. This architecture injects uniform positional information to each input contexts, ensuring consistent output generation even when the input order varies. Experiments on open domain question answering datasets, including NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and POPQA (Mallen et al., 2023), demonstrate KV-Fusion’s position-invariant nature, achieving accuracy improvements of 21.4%, 6.4%, and 6.6% over baseline models in shuffled settings. Furthermore, KV-Fusion models exhibit robust and stable accuracies even with additional contexts compared to other approaches.
@@ 2 Method
Our KV-Fusion architecture is illustrated in Figure 2. For clarity, we refer to this prefill decoder as $\mathcal{D}_{\textrm{p}}$, which is characterized by the number of key and value heads $|\mathit{H}|$, each with a dimension of $\mathit{d}_{h}$. We denote the trainable decoder as $\mathcal{D}_{\textrm{t}}$, and represent the set of input passages as $\mathcal{C}=\{c_{1},c_{2},\dots,c_{N}\}$ with fixed token length $n$ for each $c_{i}$. This set of passages represents smaller chunks of a long document or retrieved contexts. Lastly, let $L$ represent the total number of layers in $\mathcal{D}_{\textrm{p}}$ and $\mathcal{D}_{\textrm{t}}$, and let $l$ denote the $l$th layer.
@@
extracts the KV cache from multiple input passages in parallel, resulting in the injection of identical local positional embeddings ${\{\vec{p_{1}},\vec{p_{2}},\dots,\vec{p_{n}}\}}$. The layer-wise cache representation for each $c_{i}$ is as follows:
@@
Here, $K^{l},V^{l}\in\mathbb{R}^{|\mathit{H}|\times(N\times n)\times\mathit{d}_{h}}$ are reshaped KV-cache for the corresonding layer over input passages. These caches prefills and serve as grounding knowledge for training $\mathcal{D}_{\textrm{t}}$.
@@
Here, $q$ denotes the instruction with query tokens and $y$ is answer tokens. $\mathcal{C^{\prime}}$ represents the set of input passages tokens, and $\{K^{l},V^{l}\}_{l=1}^{L}$ is the reshaped KV-cache corresponding to $\mathcal{C^{\prime}}$. We illustrate the details of KV-Fusion in Appendix A.1
@@ 3.1 Datasets
We consider three open domain question answering datasets: Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and POPQA (Mallen et al., 2023) . [^1] Note that NQ and TriviaQA are filtered version from DPR (Karpukhin et al., 2020) For the base retrieval corpus, we utilize a December 2018 Wikipedia snapshot consisting of 21 million passages, following (Yen et al., 2024; Yu et al., 2024b). Lastly, we use the DPR (Karpukhin et al., 2020) as our baseline retriever to extract the top-40 passages for each dataset.  [^2] Note that we used DPR trained from scratch with its own hard negatives on December 2018 Wikipedia snapshot
@@
To enhance the robustness in RAG, we train models with irrelevant contexts (Fang et al., 2024; Yoran et al., 2024a). To this end, we draw the best gold context and extract key phrases among candidate passages by prompting gpt-4o API with a fine-grained template. If all responses are negative, the instance is discarded. Otherwise, we retain the extracted key phrases as evidence, which is later used for training. Negative contexts are sampled from DPR-retrieved passages that do not contain any answer. Each training instance consists of one gold context and 19 negative contexts. The prompt for this process and statistics of all datasets are desribed in Appendix A.2.
@@
Exact Match (EM) Accuracy is used for evaluation. (Asai et al., 2023; Mallen et al., 2023). However, we observe that as more documents are added to the input, baseline models tend to generate instrinsic knowledge or hallucinated responses Hsieh et al. (2024a). To address this, we incorporate answerability into the prompt, requiring responses to be concise, and limited to a single sentence. Lastly, we set a 48-token limit and use greedy-decoding Huang et al. (2023a) for baselines and KV-Fusions. The template and example for baseline are illustrated in Appendix A.3.
@@ 3.2 Training
Input Formating
Each input passage is formatted with ‘Title:{title}’ and ‘Context:{text}’, followed by a document boundary, ‘====’. For target tokens, we preprend a signal token, <<<|question_answering|>>>, to guide the model’s behavior during inference (Asai et al., 2023). Next, we append instruction and ‘Question:{question}’. Finally, we add answer tokens, which contain both answer string and a key phrase as evidence, as described in Section 3.1. We hypothesize that appending key phrases enhances the the model’s robustness (Thoppilan et al., 2022; Menick et al., 2022). Format examples are provided in Appendix A.4.
@@
We initialize both $\mathcal{D}_{\textrm{p}}$ and $\mathcal{D}_{\textrm{t}}$ with the Llama3-8B model (Dubey et al., 2024). We fine-tune on each dataset with a maximum learning rate of $2\times 10^{-5}$ using the AdamW. Across all dataset, we use a batch size of 64 on four A100(80G) GPUs. For the NQ and TQA datasets, models are trained for 2 epochs. For the POPQA dataset, we fine-tune it on top of TQA fine-tuned model due to its small training size. The same procedure is applied to the Llama3.1-8B. Detail hyperparameters are reported in Appendix A.5.
@@ 4 Results
Table 1: Accuracy comparison with differnet position-invariant readers. Across top-k results, KV-fusion maintains stable and the strong accuracies, while other models either degrade or exhibit relatively low accuracies.
@@
Table 2: Token-Level Match (TLM) of KV-Fusion and baseline models. While baseline models produce varying tokens when contexts are shuffled, KV-Fusion generates identical tokens, even with order perturbations.
@@
To demonstrate the position-agnostic property, we test models with the gold context placed at varying positions. For each dev dataset, we construct 10 versions by inserting gold context at every alternate location (1st, 3rd, etc.), along with an additional dev set where all 20 contexts are randomly shuffled. To manage the increased inference time, we evaluate the first 500 instances. As shown in Figure 3, KV-Llama3 maintains consistent accuracy across all datasets, regardless of the position of the gold context, while conventional Llama3 shows varying accuracy. A similar pattern is observed with KV-Llama3.1 and Llama3.1 as shown in Appendix A.6. Figure 4 emphasizes this difference: the accuracy of the baseline model drops considerably with shuffled contexts, while the KV models maintain stable performance. In the shuffled scenario, KV-Llama3 achieves higher accuracy than baselines on the NQ, TQA, and POPQA datasets, with similar trends observed for KV-Llama3.1. These findings suggest that KV-Fusion improves performance in RAG pipelines.
@@
We evaluate KV models alongside other position-agnostic methods: PAM QA (He et al., 2024), which employs multi-step reasoning to reduce position bias, and RePlug (Shi et al., 2024b), which predicts the next token based on a weighted score for each context. Across the test sets, we utilize up to 40 DPR-retrieved passages, using default settings for PAM QA and the same configurations as Llama3 for RePlug. As shown in Table 1, KV-models achieve the highest accuracy across datasets except top-5 NQ case. Notably, KV-models, originally trained with 20 passages, demonstrate strong robustness even with top-40 passages. PAM QA performs well with up to 20 passages but shows an average accuracy decline of 50.3% when scaled to the top 40. RePlug follows the similar pattern as the baselines but also experiences performance degradation. Comparable results are observed with contriever passages as shown in Appendix A.7. These results indicate that KV-Fusion enhances robustness even with large input passages within the RAG pipeline.
@@ 5 Ablations
Previous research (He et al., 2024; Hsieh et al., 2024b; Yu et al., 2024a) on positional bias primarily focuses on metrics like accuracy to assess position-agnostic behavior. It is often concluded that the model is position-agnostic, if the model maintains a consistent metric (such as accuracy). In this section, we further investigate targets token-level consistency. We define Token-Level Match (TLM) as the accuracy of Exact Match (EM) between responses in two scenarios: (1) POS1, where the gold context is at top, and (2) Shuffled, with randomly ordered contexts. Here, $p$ and $s$ denote the responses in POS1 and Shuffled settings, and $N$ is the total number of instances. We experiment with top-20 dev set in Section 4.
$TLM(p,s,N)=\frac{\sum_{i=1}^{N}EM(p_{i}=s_{i})}{N}$
The results are shown in Table 2. With greedy-decoding, both Llama3 and Llama3.1 produce tokens that vary between the two settings, resulting in low TLM scores across datasets. Specifically, their TLM scores are significantly below 20, highlighting their susceptibility to positional bias.
In constrast, KV-Fusion demonstrates nearly perfect TLM scores across all datasets. This indicates that KV-Fusion can generate identical tokens regardless of context order perturbations, highlightening its robust position-agnostic behavior and maintaining token-level consistency even in scenarios where context order is altered.
@@
To enhance the end-to-end performance of the RAG pipeline, rerankers Nogueira et al. (2020); Zhuang et al. (2023); Ma et al. (2024) are often used to prioritize relevants context while filtering out irrelevant ones (Glass et al., 2022; Yu et al., 2024b). However, we hypothesize that a robust model capable of handling irrelevant passages could eliminate the need for a reranking module. To test this, we rerank the top-40 retrieved passages from the test sets in Section 4 and retain only the top-20 passages, discarding the lower-ranked ones. We utilize point-wise rerankers: MonoT5-3BNogueira et al. (2020), Rankt5-3BZhuang et al. (2023), and RankLLama-7BMa et al. (2024). We then compared the EM accuracy of baseline models evaluated on these top-20 reranked test sets with KV-Fusion models evaluated on the original top-40 test sets.
Results are described in Table 3. Ranking and truncation improve accuracy for Llama3 across datasets. However, LLama3.1 exhibits slight gains in POPQA but a minor decline in NQ, TQA. This behavior is likely due to LLama3.1’s ability to process longer sequences, making reranking and truncation less beneficial. In contrast, KV-Fusion demonstrates consistently superior accuracy across all datasets, showing its robustness and suggesting that reranking may not be necessary for the RAG pipeline.
@@ 6 Related Works
Despite the promising capabilities of the RAG system, one major challenge is the notable drop in performance when irrelevant contexts exist during inference. (Shi et al., 2023; Oh and Thorne, 2023), along with incorrect responses even when the gold context appears in the middle (Liu et al., 2023). To address these issues, Xu et al. 2023 trained an auxiliary LLM to summarize and extract relevant contexts, while Yoran et al. 2024b proposed a simple Natural Language Inference (NLI) model to eliminate unnecessary passages. Also, He et al. 2024 suggests decomposing inference into multi-step resasoning, enabling the model to generate accurate response regardless of the context order. Other methods focus on internal features, such as adjusting position hidden states or calibrating attention biases (Hsieh et al., 2024b; Yu et al., 2024a). However, none of these approaches fully resolve a complete solution for ‘Lost in the Middle’ problem.
@@
Recent studies employing Key-Value Caches (KV caches) in the RAG pipeline have gained attention. For example, RAGCache Jin et al. (2024) leverages precomputed KV caches for retrieval and faster inference. Similarly, Cache Augmented Generation (CAG) Chan et al. (2024) augments the knowledge via KV caches, demonstrating its efficiency for long-context understanding. TurboRAG Lu et al. (2024) utilizes KV caches for training RAG pipeline, whereas our work focuses on using KV caches to address the ‘Lost in the Middle’ problem, enhancing robustness and enabling a order-invariant pipeline.

### v2

@@ Abstract
Recent advancements in Large Language Models (LLMs) underscore the necessity of Retrieval Augmented Generation (RAG) to leverage external information. However, LLMs are sensitive to the position of relevant information within contexts and tend to generate incorrect responses when such information is placed in the middle, known as the ‘Lost in the Middle’ phenomenon. In this paper, we introduce a framework that generates consistent outputs for decoder-only models, irrespective of the input context order. Experimental results for three open domain question answering tasks demonstrate position invariance, where the model is not sensitive to input context order and superior robustness to irrelevant passages compared to prevailing approaches for RAG pipelines.
@@
Philhoon Oh  Jinwoo Shin  James Thorne KAIST AI {philhoonoh,jinwoos,thorne}@kaist.ac.kr
@@ 1 Introduction
In Retrieval Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2021; Izacard et al., 2022), models first extract relevant information from a knowledge base and then incorporate this extracted information with its parametric knowledge to generate the response. This two-step approach is the de-facto approach for knowledge-intensive tasks (Lewis et al., 2021; Petroni et al., 2021).
However, decoder-only models exhibit an intrinsic positional bias, assigning more attention to tokens at the beginning or end of the input sequence while often overlooking relevant context located in the middle, a problem known as the ‘Lost in the Middle’ (Liu et al., 2023). Previous works to address this issue involve training with specific prompt (He et al., 2024) or data-intensive training (An et al., 2024). Other works aimed at modifying positional embeddings (Hsieh et al., 2024b) or reducing positional attention bias in LLMs (Yu et al., 2024a). Yet, none of these methods fully guarantee a solution to this intrinsic bias in LLMs for RAG.
@@
In this paper, we introduce a framework for decoder-only models, called Key Value Fusion (KV Fusion), to generate consistent outcomes regardless of input order as illustrated in Figure 1. KV Fusion consists of two components: a prefill decoder that extracts key-value caches in parallel and a trainable decoder that utilizes extracted key-value caches to produce consistent outcomes. This architecture injects uniform positional information into each input context, ensuring consistent output generation even when the input order varies. Experiments on open domain question answering datasets, including NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and POPQA (Mallen et al., 2023), demonstrate KV-Fusion’s position-invariant nature, achieving accuracy improvements of 21.4%, 6.4%, and 6.6% over baseline models in shuffled settings. Furthermore, KV-Fusion models exhibit robust and stable accuracies even with additional contexts compared to other approaches.
@@ 2 Method
Our KV-Fusion architecture is illustrated in Figure 2. For clarity, we refer to this prefill decoder as $\mathcal{D}_{\textrm{p}}$, characterized by the number of key and value heads $|\mathit{H}|$, each with a $\mathit{d}_{h}$ dimension. We denote the trainable decoder as $\mathcal{D}_{\textrm{t}}$, and represent the set of input passages as $\mathcal{C}=\{c_{1},c_{2},\dots,c_{N}\}$ with fixed token length $n$ for each $c_{i}$. This set of passages represents smaller chunks of a long document or retrieved contexts. Lastly, let $L$ represent the total number of layers in $\mathcal{D}_{\textrm{p}}$ and $\mathcal{D}_{\textrm{t}}$, and let $l$ denote the $l$th layer.
@@
extracts the KV-caches from multiple input passages in parallel, resulting in the injection of identical local positional embeddings ${\{\vec{p_{1}},\vec{p_{2}},\dots,\vec{p_{n}}\}}$. The layer-wise cache representation for each $c_{i}$ is as follows:
@@
Here, $K^{l},V^{l}\in\mathbb{R}^{|\mathit{H}|\times(N\times n)\times\mathit{d}_{h}}$ are reshaped KV-cache for the corresponding layer over input passages. These caches prefill and serve as grounding knowledge for training $\mathcal{D}_{\textrm{t}}$.
@@
Here, $q$ denotes the instruction with query tokens, and $y$ is answer tokens. $\mathcal{C^{\prime}}$ represents the set of input passages tokens, and $\{K^{l},V^{l}\}_{l=1}^{L}$ is the reshaped KV-caches corresponding to $\mathcal{C^{\prime}}$. We illustrate the details of KV-Fusion in Appendix A.1
@@ 3.1 Datasets
We consider three open domain question answering datasets: Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and POPQA (Mallen et al., 2023) . [^1] Note that NQ and TriviaQA are filtered version from DPR (Karpukhin et al., 2020) For the base retrieval corpus, we utilize a December 2018 Wikipedia snapshot consisting of 21 million passages, following (Yen et al., 2024; Yu et al., 2024b). Lastly, we use the DPR (Karpukhin et al., 2020) as our baseline retriever to extract the top 40 passages for each dataset.  [^2] Note that we used DPR trained from scratch with its hard negatives on the December 2018 Wikipedia snapshot
@@
To enhance the robustness in RAG, we train models with irrelevant contexts (Fang et al., 2024; Yoran et al., 2024a). To this end, we draw the best gold context and extract key phrases among candidate passages by prompting GPT-4o API with a fine-grained template. If all responses are negative, the instance is discarded. Otherwise, we retain the extracted key phrases as evidence, which is later used for training. Negative contexts are sampled from DPR-retrieved passages that do not contain any answer. Each training instance consists of one gold context and 19 negative contexts. The prompt for this process and statistics of all datasets are described in Appendix A.2.
@@
Exact Match (EM) Accuracy is used for evaluation. (Asai et al., 2023; Mallen et al., 2023). However, we observe that as more documents are added to the input, baseline models tend to generate intrinsic knowledge or hallucinated responses Hsieh et al. (2024a). To address this, we incorporate answerability into the prompt, requiring responses to be concise and limited to a single sentence. Lastly, we set a 48-token limit and use greedy-decoding Huang et al. (2023a) for baselines and KV-Fusions. The template and example for baseline are illustrated in Appendix A.3.
@@ 3.2 Training
Input Formatting
Each input passage is formatted with ‘Title:{title}’ and ‘Context:{text}’, followed by a document boundary, ‘====’. For target tokens, we preprend a signal token, <<<|question_answering|>>>, to guide the model’s behavior during inference (Asai et al., 2023). Next, we append instruction and ‘Question:{question}’. Finally, we add answer tokens, which contain both an answer string and a key phrase as evidence, as described in Section 3.1. We hypothesize that appending key phrases enhances the model’s robustness (Thoppilan et al., 2022; Menick et al., 2022). Format examples are provided in Appendix A.4.
@@
We initialize both $\mathcal{D}_{\textrm{p}}$ and $\mathcal{D}_{\textrm{t}}$ with the Llama3-8B model (Dubey et al., 2024). We fine-tune each dataset with a maximum learning rate of $2\times 10^{-5}$ using the AdamW. Across all datasets, we use a batch size of 64 on four A100(80G) GPUs. For the NQ and TQA datasets, models are trained for two epochs. For the POPQA dataset, we fine-tune it on top of the TQA fine-tuned model due to its small training size. The same procedure is applied to the Llama3.1-8B. Detail hyperparameters are reported in Appendix A.5.
@@ 4 Results
Table 1: Percentage accuracy comparison between position-invariant readers. Across top-k results, KV-fusion maintains stable and the strong accuracies, while other models either degrade or exhibit relatively low accuracies.
@@
Table 2: Percentage Token-Level Match of KV-Fusion and baseline models. While baselines produce varying tokens when contexts are shuffled, KV-Fusion generates identical tokens, even with order perturbations.
@@
To demonstrate the position-agnostic property, we test models with the gold context placed at varying positions. For each dev dataset, we construct 10 versions by inserting gold context at every alternate location (1st, 3rd, etc.), along with an additional dev set where all 20 contexts are randomly shuffled. To manage the increased inference time, we evaluate the first 500 instances. As shown in Figure 3, KV-Llama3 maintains consistent accuracy across all datasets, regardless of the position of the gold context, while conventional Llama3 shows varying accuracy. A similar pattern is observed with KV-Llama3.1 and Llama3.1 as shown in Appendix A.6. Figure 4 emphasizes this difference: the accuracy of the baseline model drops considerably with shuffled contexts, while the KV models maintain stable performance. In the shuffled scenario, KV-Llama3 achieves higher accuracy than baselines on the NQ, TQA, and POPQA datasets, with similar trends observed for KV-Llama3.1. These findings suggest that KV-Fusion improves performance in RAG.
@@
We evaluate KV models alongside other position-agnostic methods: PAM QA (He et al., 2024), which employs multi-step reasoning to reduce position bias, and RePlug (Shi et al., 2024b), which predicts the next token based on a weighted score for each context. Across the test sets, we utilize up to 40 DPR-retrieved passages, using default settings for PAM QA and the same configurations as Llama3 for RePlug. As shown in Table 1, KV-models achieve the highest accuracy across datasets except the top-5 NQ case. Notably, KV models, originally trained with 20 passages, demonstrate strong robustness even with the top 40 passages. PAM QA performs well with up to 20 passages but shows an average accuracy decline of 50.3% when scaled to the top 40. RePlug follows a similar pattern as the baselines but also experiences performance degradation. Comparable results are observed with contriever passages as shown in Appendix A.7. These results indicate that KV-Fusion enhances robustness even with large input passages within the RAG pipeline.
@@ 5 Ablations
Previous research (He et al., 2024; Hsieh et al., 2024b; Yu et al., 2024a) on positional bias study the effect of performance (e.g. accuracy) under perturbation of the input. Models are regarded as consistent if performance is stable regardless of the position of the gold context. In this section, we further investigate target token-level consistency. We define Token-Level Match (TLM) as the accuracy of Exact Match (EM) between responses in two scenarios: (1) POS1, where the gold context is at the top, and (2) Shuffled, with randomly ordered contexts. Here, $p$ and $s$ denote the responses in POS1 and Shuffled settings, and $N$ is the total number of instances. We experiment with the top-20 dev set in Section 4.
$TLM(p,s,N)=\frac{1}{N}{\sum_{i=1}^{N}EM(p_{i}=s_{i})}$
The results are shown in Table 2. Using greedy-decoding (as this is non-stochastic), Llama3 and Llama3.1 are sensitive to positional context, resulting in low TLM scores across datasets. Notably, the models are most sensitive to positional bias on POPQA. In contrast, KV-Fusion demonstrates nearly perfect TLM scores across all datasets (all $>99\%$). This indicates that KV-Fusion generates near identical tokens regardless of context-order perturbations, highlighting position-agnostic behavior and maintaining token-level consistency where context order is perturbed.
@@
To enhance the end-to-end performance of the RAG pipeline, rerankers Nogueira et al. (2020); Zhuang et al. (2023); Ma et al. (2024) are often used to prioritize relevant context while filtering out irrelevant ones (Glass et al., 2022; Yu et al., 2024b). However, we hypothesize that a robust model capable of handling irrelevant passages could eliminate the need for a reranking module. To test this, we rerank the top 40 retrieved passages from the test sets in Section 4 and retain only the top 20 passages, discarding the lower-ranked ones. We utilize point-wise rerankers: MonoT5-3BNogueira et al. (2020), Rankt5-3BZhuang et al. (2023), and RankLLama-7BMa et al. (2024). We then compared the EM accuracy of baseline models evaluated on these top-20 reranked test sets with KV-Fusion models evaluated on the original top-40 test sets.
Results are described in Table 3. Ranking and truncation improve the accuracy of Llama3 across datasets. However, LLama3.1 exhibits slight gains in POPQA but a minor decline in NQ, TQA. This behavior is likely due to LLama3.1’s ability to process longer sequences, making reranking and truncation less beneficial. In contrast, KV-Fusion demonstrates consistently superior accuracy across all datasets, showing its robustness and suggesting that reranking may not be necessary for the RAG.
@@ 6 Related Works
Despite the promising capabilities of the RAG system, one major challenge is the notable drop in performance when irrelevant contexts exist during inference. (Shi et al., 2023; Oh and Thorne, 2023), along with incorrect responses even when the gold context appears in the middle (Liu et al., 2023). To address these issues, Xu et al. 2023 trained an auxiliary LLM to summarize and extract relevant contexts, while Yoran et al. 2024b proposed a simple Natural Language Inference (NLI) model to eliminate unnecessary passages. Also, He et al. 2024 suggests decomposing inference into multi-step reasoning, enabling the model to generate accurate responses regardless of the context order. Other methods focus on internal features, such as adjusting position hidden states or calibrating attention biases (Hsieh et al., 2024b; Yu et al., 2024a). However, none of these approaches fully resolve a complete solution for the ‘Lost in the Middle’ problem.
@@
Recent studies in Key-Value Cache (KV cache) in RAG focus on improving long context understanding and runtime efficiency. For example, RAGCache Jin et al. (2024) leverages precomputed KV caches for retrieval and faster inference. Similarly, Cache Augmented Generation (CAG) Chan et al. (2024) augments the knowledge via KV caches, demonstrating its efficiency for long-context understanding. TurboRAG Lu et al. (2024) utilizes KV caches for training the RAG pipeline, whereas our work focuses on using KV caches to address the ‘Lost in the Middle’ problem, enhancing robustness and enabling an order-invariant pipeline.